A compound was found to contain 29.1% Na, 40.5% S, and 30.4% O. Find the empirical formula for this compound
1
Expert's answer
2018-04-08T08:16:10-0400
.ω (Na) =29.1 %
.ω(S) =40.5 %
.ω (O)=30.4 %
In 100 g of compound there are 29.1 g of Na, 40.5 g of S and 30.4 g of O
Find the amount of substance of each element
.n(Na)= (m(Na))/(A_r (Na)) =29.1/23 =1.27 mol
.n(S) = (m(S))/(Ar(S))=40.5/32 =1.27 mol
.n(O)= (m(O))/(A_r (O)) = 30.4/16 =1.9 mol
.n(Na):n(S):n(O)=1.27:1.27:1.9=1:1:1.49
Multiply the result by 2 and have n(Na):n(S):n(O)=2:2:3
The formula is Na_2 S_2 O_3- sodium thiosulfate